# Task-Unaware Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

### v1

@@ 6 Conclusion and Discussion
In this paper, we introduced a novel task-unaware lifelong robot learning framework that combines retrieval-based local adaptation with selective weighting during test time. Our approach enables robots to continuously learn and adapt in dynamic environments without explicit task identifiers or predefined boundaries. Leveraging an episodic memory $\mathcal{M}$, our method retrieves relevant past demonstrations based on visual and language similarities, allowing the robot to fine-tune its policy locally. The selective weighting mechanism enhances adaptation by prioritizing the most challenging segments of the retrieved demonstrations. Notably, our framework is not only robust, but is compatible with various memory-based lifelong learning methods, enhancing a robot’s ability to perform previously learned tasks as a paradigm.
A key challenge lies in the selective weighting process, particularly in finding the Separation Segment. Real-world noise, the multimodal nature of manipulation actions, and varying semantic information make it sometimes difficult to accurately identify Separation Segment in demonstration trajectories. Addressing this issue will be the focus of our future work to further improve our approach.
@@ A.2.2 Training Hyperparameters
Table 6 provides a summary of the essential hyperparameters used during training and local adaptation. The model was trained on a combination of A40, A100, and L40S GPUs, while we also leveraged multi-GPU configurations to accelerate the training process. For each task, demonstration data was initially collected and provided by LIBERO benchmark. However, due to version discrepancies that introduced visual and physical variations in the simulation, we reran the demonstrations with the latest version to obtain updated observations. It is important to note that occasional rollout failures occurred because different versions of RoboMimic Simulation (Mandlekar et al., 2021) utilize varying versions of the MuJoCo Engine (Todorov et al., 2012).
@@
Number of Demos per Task in EM $8$
@@ A.2.3 Baseline Details
We follow the implementation of all baselines and hyperparameters for individual algorithms from (Liu et al., 2024), maintaining the same backbone model and episodic memory structure as in our approach. During the training phase, we also apply the same learning hyperparameters outlined in Table 6.
A.3 Details about Task-unaware setting
@@ A.3.1 Datasets Structure
Our dataset inherent the dataset from LIBERO (Liu et al., 2024), maintaining all the attributes and data. Additionally, we add demo description to each demonstration to achieve task unawareness and augmented description to augment language description during training (See Figure 5). Unlike the dataset from LIBERO, which groups demonstrations together under one specific task, our dataset wrap all demonstrations with random order to eliminate the task boundary.
@@ A.3.2 Description Paraphrase
We leverage the Phi-3-mini-4k-instruct model (mini-4k instruct, 2024) to paraphrase the task description. The process and prompts that we use are illustrated in Figure 6.
@@ A.4.1 Detailed Heuristics and Implementations
To assign weights to retrieved demonstrations, we analyze the distance between demonstration and failed rollout trajectories. Typically, the comparison distance increases as the failed rollout diverges from the demonstration.
@@
Due to the multi-modal nature of robotic actions and visual observation noise, raw distance comparisons can be erratic. To address this, we smooth the distance curves using a moving window. Despite smoothing, the trend may remain jittery, making it challenging to pinpoint a single separation point where performance deviates. Instead, we identify a range of frames representing the Separation Segment where the distances worsen, indicating vulnerable steps in a manipulation task.
@@
We apply two thresholds to detect the segment. Specifically, we locate frames where the distance falls between $\frac{1}{8}$ and $\frac{1}{3}$ of the maximum observed distance. We focus on the last occurrence within this range to account for possible initial divergent paths that later converge. Once identified, we extend the separation segment by $15$ frames before and after to mitigate noise effects.
@@
For each frame within the separation segment, we add a weight of $0.3$ to the initially uniform weight vector. This process is repeated for up to five failed rollouts per retrieved demonstration. After processing all demonstrations, we clip the weights to a maximum of $2$ and normalize the weight vector to maintain a consistent loss function scale, ensuring stable gradient updates.
During adaptation, the resulting weights ($w_{t,n}$) are integrated into the loss function as described in  equation 3. This approach enhances the influence of critical samples while reducing the impact of less relevant ones, thereby improving the model’s learning efficiency.
A.4.2 Detailed Ablation Studies on Selective Weighting.
The average success rate per benchmark is illustrated in Table 2. The detailed results on each task are shown in Table 7, Table 8, and Table 9.
Table 7: Ablation Study Results on libero_object: Average Success Rates and Standard Devision for Each Task Across Epochs.
Method Task  Epoch ULA WLA
Epoch 15 Epoch 20 Epoch 25 Epoch 15 Epoch 20 Epoch 25
0 0.68 $\pm$ 0.04 0.37 $\pm$ 0.12 0.62 $\pm$ 0.04 0.57 $\pm$ 0.14 0.67 $\pm$ 0.15 0.57 $\pm$ 0.04
1 0.20 $\pm$ 0.08 0.40 $\pm$ 0.13 0.35 $\pm$ 0.22 0.35 $\pm$ 0.15 0.45 $\pm$ 0.06 0.13 $\pm$ 0.06
2 0.77 $\pm$ 0.14 0.85 $\pm$ 0.06 0.78 $\pm$ 0.15 0.90 $\pm$ 0.08 0.78 $\pm$ 0.14 0.82 $\pm$ 0.11
3 0.68 $\pm$ 0.15 0.78 $\pm$ 0.06 0.70 $\pm$ 0.03 0.70 $\pm$ 0.09 0.60 $\pm$ 0.10 0.75 $\pm$ 0.08
4 0.75 $\pm$ 0.08 0.87 $\pm$ 0.02 0.78 $\pm$ 0.07 0.70 $\pm$ 0.06 0.78 $\pm$ 0.07 0.88 $\pm$ 0.03
5 0.47 $\pm$ 0.19 0.65 $\pm$ 0.05 0.53 $\pm$ 0.04 0.37 $\pm$ 0.09 0.42 $\pm$ 0.07 0.60 $\pm$ 0.13
6 0.52 $\pm$ 0.06 0.53 $\pm$ 0.09 0.38 $\pm$ 0.16 0.65 $\pm$ 0.12 0.52 $\pm$ 0.12 0.55 $\pm$ 0.08
7 0.47 $\pm$ 0.19 0.58 $\pm$ 0.14 0.57 $\pm$ 0.09 0.58 $\pm$ 0.04 0.73 $\pm$ 0.04 0.60 $\pm$ 0.18
8 0.55 $\pm$ 0.10 0.58 $\pm$ 0.17 0.50 $\pm$ 0.13 0.58 $\pm$ 0.06 0.70 $\pm$ 0.09 0.72 $\pm$ 0.09
9 0.70 $\pm$ 0.18 0.45 $\pm$ 0.10 0.58 $\pm$ 0.03 0.40 $\pm$ 0.15 0.58 $\pm$ 0.02 0.53 $\pm$ 0.09
Table 8: Ablation Study Results on libero_goal: Average Success Rates and Standard Devision for Each Task Across Epochs.
Method Task  Epoch ULA WLA
Epoch 15 Epoch 20 Epoch 25 Epoch 15 Epoch 20 Epoch 25
0 0.62 $\pm$ 0.09 0.75 $\pm$ 0.05 0.68 $\pm$ 0.10 0.72 $\pm$ 0.04 0.83 $\pm$ 0.09 0.65 $\pm$ 0.03
1 0.88 $\pm$ 0.03 0.92 $\pm$ 0.03 0.88 $\pm$ 0.02 0.87 $\pm$ 0.06 0.88 $\pm$ 0.04 0.92 $\pm$ 0.04
2 0.65 $\pm$ 0.13 0.72 $\pm$ 0.12 0.80 $\pm$ 0.03 0.68 $\pm$ 0.08 0.70 $\pm$ 0.08 0.83 $\pm$ 0.06
3 0.38 $\pm$ 0.07 0.25 $\pm$ 0.03 0.32 $\pm$ 0.09 0.32 $\pm$ 0.12 0.38 $\pm$ 0.16 0.32 $\pm$ 0.06
4 0.88 $\pm$ 0.04 0.80 $\pm$ 0.05 0.82 $\pm$ 0.03 0.87 $\pm$ 0.04 0.77 $\pm$ 0.14 0.92 $\pm$ 0.04
5 0.60 $\pm$ 0.10 0.53 $\pm$ 0.13 0.63 $\pm$ 0.20 0.58 $\pm$ 0.07 0.62 $\pm$ 0.12 0.77 $\pm$ 0.03
6 0.15 $\pm$ 0.03 0.15 $\pm$ 0.09 0.22 $\pm$ 0.07 0.13 $\pm$ 0.04 0.22 $\pm$ 0.06 0.20 $\pm$ 0.03
7 0.93 $\pm$ 0.04 0.95 $\pm$ 0.05 1.00 $\pm$ 0.00 0.97 $\pm$ 0.02 0.93 $\pm$ 0.02 0.93 $\pm$ 0.04
8 0.78 $\pm$ 0.07 0.80 $\pm$ 0.03 0.77 $\pm$ 0.04 0.72 $\pm$ 0.06 0.67 $\pm$ 0.11 0.90 $\pm$ 0.05
9 0.25 $\pm$ 0.03 0.33 $\pm$ 0.08 0.50 $\pm$ 0.10 0.43 $\pm$ 0.17 0.23 $\pm$ 0.06 0.32 $\pm$ 0.09
Table 9: Ablation Study Results on libero_spatial: Average Success Rates and Standard Devision for Each Task Across Epochs.
Method Task  Epoch ULA WLA
Epoch 15 Epoch 20 Epoch 25 Epoch 15 Epoch 20 Epoch 25
0 0.35 $\pm$ 0.18 0.33 $\pm$ 0.07 0.47 $\pm$ 0.07 0.45 $\pm$ 0.10 0.45 $\pm$ 0.10 0.42 $\pm$ 0.13
1 0.48 $\pm$ 0.09 0.43 $\pm$ 0.04 0.48 $\pm$ 0.10 0.30 $\pm$ 0.05 0.58 $\pm$ 0.19 0.40 $\pm$ 0.13
2 0.32 $\pm$ 0.11 0.35 $\pm$ 0.13 0.28 $\pm$ 0.12 0.40 $\pm$ 0.19 0.50 $\pm$ 0.13 0.45 $\pm$ 0.13
3 0.48 $\pm$ 0.03 0.47 $\pm$ 0.09 0.60 $\pm$ 0.05 0.48 $\pm$ 0.07 0.47 $\pm$ 0.11 0.50 $\pm$ 0.00
4 0.17 $\pm$ 0.04 0.30 $\pm$ 0.03 0.13 $\pm$ 0.07 0.22 $\pm$ 0.07 0.18 $\pm$ 0.07 0.23 $\pm$ 0.02
5 0.12 $\pm$ 0.09 0.20 $\pm$ 0.08 0.28 $\pm$ 0.09 0.25 $\pm$ 0.10 0.22 $\pm$ 0.09 0.27 $\pm$ 0.02
6 0.60 $\pm$ 0.13 0.58 $\pm$ 0.02 0.47 $\pm$ 0.10 0.57 $\pm$ 0.07 0.58 $\pm$ 0.07 0.67 $\pm$ 0.03
7 0.52 $\pm$ 0.06 0.42 $\pm$ 0.02 0.38 $\pm$ 0.07 0.38 $\pm$ 0.06 0.38 $\pm$ 0.04 0.38 $\pm$ 0.06
8 0.30 $\pm$ 0.05 0.42 $\pm$ 0.08 0.30 $\pm$ 0.00 0.40 $\pm$ 0.10 0.28 $\pm$ 0.03 0.30 $\pm$ 0.03
9 0.20 $\pm$ 0.10 0.32 $\pm$ 0.09 0.42 $\pm$ 0.03 0.17 $\pm$ 0.07 0.33 $\pm$ 0.06 0.17 $\pm$ 0.04
@@ A.5 Detailed Testing Results
We selected 20 typical scenarios among libero_90. The list of those scenarios can be found in Table 10. Additionally, the testing results of our method and baselines including ER-WLA, ER, Packnet, are listed in Table 11
@@ 20 Pick up the book in the middle and place it on the cabinet shelf Study scene4
Table 11: Detailed Comparisons on libero_different_scenes Benchmark. It illustrates that after the reaching the capacity of PackNet, it could no longer deal with new tasks anymore.
Task ER-WLA ER Packnet

### v3

@@ 6 Conclusion and Discussion
In this paper, we introduced a novel task-free lifelong robot learning framework that combines retrieval-based local adaptation with selective weighting before policy deployment. Our approach enables robots to continuously learn and restore forgotten knowledge and skills in dynamic environments without relying on task boundaries or IDs. Notably, our framework is not only robust, but is compatible with various memory-based lifelong learning methods, enhancing a robot’s ability to perform previously learned tasks as a paradigm.
A limitation of our framework is the scalability of the storage memory $\mathcal{M}$, as we continuously accumulate demonstrations. However, since image embeddings—serving dual purposes (input to the manipulation policy and data retrieval for local adaptation)—are generated by a pre-trained model, our approach is naturally extendable: this allows for significant storage reduction in future implementations, by simply storing smaller embeddings instead of raw images in $\mathcal{M}$.
@@ A.1 Notations
Algorithm 1 RWLA for Task-free Lifelong Robot Learning
Lifelong Learning Phase:
  1. 1.
    Initialize model parameter $\theta$, storage memory ${}\mathcal{M}=\{\}$, and tasks $\{\mathcal{T}_{i}\}$, $i=1,2,\dots,T$
  2. 2.
    $K\in\{1,2,\dots,T\}$
    1. (a)
      Train $\theta$ on $\mathcal{D}_{K}\cup\mathcal{M}$ using Eq 1
    2. (b)
      Randomly store a small number of demonstrations from $\mathcal{D}_{K}$ into $\mathcal{M}$
During deployment, robot encounters a testing scenario $\mathcal{S}_{deploy}\sim p(\mathcal{T}_{i}),1\leq i\leq T$:
Reviewing Phase:
  1. 1.
    Rollout $10$ episodes on $\mathcal{S}_{deploy}$ to assess robot’s performance with $\theta$
  2. 2.
    Retrieve $\tilde{N}$ demonstrations from $\mathcal{M}$ based on embedding distance using Eq 4.1 (4.1)
  3. 3.
    Compute $w_{t,n}$ based on selective weighting (4.2.1)
  4. 4.
    $\theta^{\prime}\leftarrow$ Locally adapt $\theta$ using Eq 2 as skill restoration within limited epochs (4.2.2)
Testing Phase: Test $\theta^{\prime}$ in $\mathcal{S}_{deploy}$
@@ A.2.2 Training Hyperparameters
Table 6 provides a summary of the essential hyperparameters used during training and local adaptation. The model training was conducted using a combination of A40, A100, and L40S GPUs in a multi-GPU configuration to optimize the training process. This distributed computing setup significantly enhanced efficiency, reducing the training time per benchmark from 12 hours on a single GPU to 6 hours using 3 GPUs in parallel. For each task, demonstration data was initially collected and provided by LIBERO benchmark. However, due to version discrepancies that introduced visual and physical variations in the simulation, we reran the demonstrations with the latest version to obtain updated observations. It is important to note that occasional rollout failures occurred because different versions of RoboMimic Simulation (Mandlekar et al., 2021) utilize varying versions of the MuJoCo Engine (Todorov et al., 2012).
@@
Number of Demos per Task in $\mathcal{M}$ ($\tilde{N}$) $8$
@@ A.2.3 Baseline Details
We follow the implementation of baselines and hyperparameters for individual algorithms from (Liu et al., 2024), maintaining the same backbone model and storage memory structure as in our approach. During training, we also apply the same learning hyperparameters outlined in Table 6.
A.3 Details about Blurred Task Boundary setting
@@ A.3.1 Datasets Structure
Our dataset inherent the dataset from LIBERO (Liu et al., 2024), maintaining all the attributes and data. Additionally, we add demo description to each demonstration to blur task boundary and augment language description during training (See Figure 6). Unlike the dataset from LIBERO, which groups demonstrations together under one specific task, our dataset wrap all demonstrations with random order to eliminate the task boundary.
@@ A.3.2 Description Paraphrase
We leverage the Phi-3-mini-4k-instruct model (mini-4k instruct, 2024) to paraphrase the task description. The process and prompts that we use are illustrated in Figure 7. As shown for the libero_spatial task in Figure 8, both BERT and Sentence Similarity Model struggle to distinguish tasks based on embeddings from the paraphrased descriptions. This observation further underscores the task-blurry setting in our experiments.
@@
(a) PCA Visualization based on BERT
(b) PCA Visualization based on SS Model
Figure 8: Task Blurry Effect on libero_spatial benchmark. After paraphrasing the task descriptions, both BERT and SS models struggle to distinguish the tasks in libero_spatial.
Figure 9: Hyperparameter Sensitivity Check.
@@ A.4.1 Detailed Heuristics and Implementations
To assign weights to retrieved demonstrations, we analyze the image embedding distance between demonstration and failed rollout trajectories. Typically, the embedding distance increases as the failed rollout diverges from the demonstration. We selectively add weights for the frames in the retrieved demonstration using the Embedding Distance Curve (EDC), derived from the Embedding Distance Matrix (EDM), as illustrated in Figure 10.
@@
Due to the multi-modal nature of robotic actions and visual observation noise, raw embedding distances can be erratic. To mitigate this, we smooth the EDC using a moving average window. Despite smoothing, the curve may remain jittery, making it difficult to pinpoint a single divergence point. Therefore, we identify a range of frames, termed the Separation Segment, where the distances increase, indicating vulnerable steps that lead to task failure.
@@
We apply two thresholds to identify the segment: a lower threshold at $\frac{1}{8}$ and an upper threshold at $\frac{1}{3}$ of the maximum observed distance in EDC. We locate frames where the smoothed embedding distance falls within this range, focusing on the last occurrence to account for initial divergences that may later converge. We then extend this segment by $15$ frames before and after to mitigate noise effects.
@@
For each frame within the Separation Segment, we increment the corresponding weight in the initially uniform weight vector by $0.3$. This process is repeated for up to five failed rollouts per retrieved demonstration. After processing all demonstrations, we clip the weights to a maximum of $2$ and normalize the weight vector to maintain consistent loss scaling and ensure stable gradient updates. During adaptation, the resulting weights ($w_{t,n}$) are integrated into the loss function as described in Equation (2). This selective weighting emphasizes critical samples while reducing the influence of less relevant ones, thereby enhancing the model’s learning efficiency.
A.4.2 Hyperparameter Sensitivity Analysis on Selective Weighting.
Figure 9 presents the sensitivity analysis of the hyperparameters—lower threshold ($\frac{1}{8}$), higher threshold ($\frac{1}{3}$), and padding step ($15$ steps)—used to identify Separation Segments during selective weighting. The experiments are conducted on three random seeds as well. The results demonstrate that our proposed method’s performance is robust to variations in these hyperparameters.
@@ A.5 Detailed Testing Results
We selected 20 typical scenarios among libero_90. The list of those scenarios can be found in Table 7. Additionally, the testing results of our method and baselines including ER-RWLA, ER, Packnet, are listed in Table 8
A.6 Discussion on Potential Forgetting during Local Adaptation
Our method addresses this issue through a robust deployment strategy. After sequential learning, we preserve the final model as a stable foundation. For each testing scenario, we fine-tune a copy of this model using our weighted local adaptation mechanism. Crucially, we always return to the preserved final model for subsequent scenarios, ensuring that each adaptation starts from the same well-trained baseline and previous adaptations do not influence future ones. This approach keeps local adaptations isolated and prevents the accumulation of forgetting effects.
@@ 20 Pick up the book in the middle and place it on the cabinet shelf Study scene4
Table 8: Detailed Comparisons on libero_different_scenes Benchmark. It illustrates that after reaching the capacity of PackNet, its performance on new tasks would drop drastically.
Task ER-RWLA ER Packnet
@@
Figure 10: Illustration of the selective weighting heuristic using (a) Embedding Distance Matrix (EDM) and (b) Embedding Distance Curve (EDC). In the demonstration, the robot successfully picks up a jar and places it into a basket. In the failed rollout, the robot fails during the picking stage, resulting in the absence of subsequent steps. The steps surrounding the picking procedure are identified as the Separation Segment and are assigned higher weights during adaptation to address the model’s shortcomings. Specifically, the Separation Segment is determined by the smoothed minimum $L_{2}$ distances from EDC—obtained from EDM, where each of its entry indicates the embedding distance between a demonstration and failed rollout frame, as shown in this figure.